2. Assume demand of a ladies fashion dress is normally distributed with mean µ = 3,192 and standard deviation  = 1,181. Selling Price = $180, Purchase price = $110, Salvage value = $90.

(a) Compute the target cycle service level.

(b) Find the z value for the target cycle service level of part (a).

(c) Find the optimal order quantity.
Answer #1

Answer: - According to given data

a) cycle service level formula = Cu/Cu+Co

Cu = underage cost = selling price - cost price = $180 - $110 = $70

Co = Overage cost = Cost price - salvage cost = $110 - $90 = $20

Cycle service level = $70 / $70+$20 = $70 / $90 = 0.7778

b) Z value for the obtained service cycle is obtained by standard probability table

Z value = 0.76

c) Formula to calculate the order quantity = mue + z * standard deviation

Optimal order quantity = 3192 + 0.76*1181 = 4089.56 ~ 4090